OpenStudy (anonymous):

if a rubik's cube has a volume of 384 cubic centimeters, how long is one side of the cube? (recall that the volume of a cube is calculated by L3, where L is the length of one side.

OpenStudy (anonymous):

A cube has all sides equal so sqrt384 twice.

OpenStudy (anonymous):

Let l be side of the cube.. => its volume = l^3 here, 384 = l^3 => l = cube root of 384

OpenStudy (nubeer):

let x be the side vol =x*x*x x^3 = 384 \[x = \sqrt[3]{384}\]